Compute fxy and fyx for the function f(x,y) = e^(x+y) sin x + 9x^2 + 2xy at (1,2)
1
Expert's answer
2020-04-14T16:42:59-0400

"f_x'=e^{x+y}\sin x+e^{x+y}\cos x+18x+2y" , "f_{xy}''=e^{x+y}\sin x+e^{x+y}\cos x+2"

"f_y'=e^{x+y}\sin x+2x" , "f_{yx}''=e^{x+y}\sin x+e^{x+y}\cos x+2" then

"f_{xy}''=f_{yx}''=e^{x+y}(\sin x+\cos x)+2\approx29.7" at "(1,2)"